OpenStudy (anonymous):

Please help terrible at graphs What is the graph of the inequality? -3y>=6x - 3

OpenStudy (anonymous):

@1on1

OpenStudy (anonymous):

@JakeV8

OpenStudy (anonymous):

solve for the y value first, before you try to graph. you have that "-3" in front of the y, so you need to divide both sides by -3 to get y all alone. However, watch out!! :) This is an inequality, and one of the rules is that if you multiply or divide both sides by a negative number, you have to switch the direction of the inequality to point the other way. So, -3y ≥ 6x - 3 -3y / (-3) ≤ (6x - 3) / (-3)

OpenStudy (anonymous):

uh ok

OpenStudy (anonymous):

i seriously have no idea what to do i dont understand how you got all that stuff after so,

OpenStudy (anonymous):

lik e i dont get the second line

OpenStudy (anonymous):

-3y / (-3) ≤ (6x - 3) / (-3) ?????

OpenStudy (tyteen4a03):

\(\frac{-3y}{-3} ≤ \frac{6x-3}{-3}\).

OpenStudy (tyteen4a03):

You're dividing both sides by the same value so you get the inequality with the subject y.

OpenStudy (anonymous):

is the shaded part on the uper right of the graph?

OpenStudy (anonymous):



OpenStudy (anonymous):

?

OpenStudy (tyteen4a03):

No. It's the region below the line.

OpenStudy (anonymous):

what????

OpenStudy (aripotta):

you could always put the inequality in geogebra to get your answer

OpenStudy (aripotta):

but if you have to explain it, then don't

OpenStudy (anonymous):



OpenStudy (anonymous):

ive been on this question for an hour this is annoying why can i cant just get an answer i guess people forget im taking timed tests

OpenStudy (anonymous):

plus i have no freakin idea what any of this is

OpenStudy (anonymous):

\[y \le 1-2x is the formula graph one \sec\]

OpenStudy (anonymous):

what is that crap

OpenStudy (anonymous):

\[y \le 1-2x\]

OpenStudy (anonymous):

that is the formula and graphing give me a sec

OpenStudy (aripotta):

use geogebra for graphs so you don't waste your time

OpenStudy (anonymous):

yeah what ari potta posted is the graph

OpenStudy (anonymous):

I like wolframalpha a little better

OpenStudy (anonymous):

at least someone answered my question after an hour thank you @AriPotta

OpenStudy (aripotta):

i've heard of wolfram, but personally haven't used it before